Find the equation of the line through the given point that is perpendicular to the given line. y=-2/3x+7, R(4,-1)
Answer by jmg(22)

In the equation y = -2/3x + 7 , I am assuming the x is not on the bottom with the 3 ??????? If it is you can't work this problem in this manner. As long as the problem is (2/3)x, this is how you work it.
In the given equation the slope is the number in front of x. So the slope is
-2/3.
A line perpendicular to that line would have the negative reciprocal slope.
So the slope of the perpendicular line would be 3/2 (flip it and change the sign to the opposite)
Then you fill into point-slope form:
y - y sub 1 = m (x - x sub 1), where m is the slope and x sub 1 and y sub 1 are the coordinates of point R(4, -1) {{ sub means subscript by the way}}
Fill in the formula
y - (-1) = 3/2 (x - 4)
y + 1 = (3/2)x - 6
y = (3/2)x - 7
This is the answer in slope-intercept form. You did not specify what form to write it in!